# Learning High-Order Relationships of Brain Regions

### 02 Dec 2023 v1

@@ 4.2 Optimization Framework
We propose a new IB framework named multi-head drop-bottleneck to optimize HyBRiD. To adopt an information bottleneck view of HyBRiD, we consider $X$, $Y$ and $H$ are random variables in the Markovian chain $X\leftrightarrow Y\leftrightarrow H$. According to our MIMR objective, we optimize
$\arg\max I(H;Y)-\beta I(H;X)$ (8)
where $\beta$ is a coefficient trading off informativeness and redundancy. Since optimizing the mutual information for high-dimensional continuous variables is intractable, we instead optimize the lower bound of Equation 8. Specifically, for the first term (informativeness), it is easy to show
@@
$=\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]+\mathbb{E}_{p(H)}[\operatorname{KL}(p(Y|H)|q_{\phi}(Y|H))]$
$\geq\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]$
where $\mathbb{H}[\cdot]$ is the entropy computation. Since there is no learnable component in the entropy of $Y$, we only need to optimize the second term $\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]$. $q_{\phi}$ can be considered as a model that predicts $Y$ based on $H$, which essentially corresponds to $\mathcal{F}_{l}\circ\mathcal{F}_{w}$, where $\circ$ is the function composition. In practice, we set $q_{\phi}$ as a Gaussian model with variance $1$ most probabilistic machine learning models for continuous data modeling. (Alemi et al., 2016; Luo et al., 2019; Peng et al., 2018; Kingma & Welling, 2013).
@@
$I(H;X)\leq\sum_{k=1}^{K}I(\bm{h}^{k};X)\leq\sum_{k=1}^{K}\sum_{i=1}^{N}I(\bm{h}^{k}_{i};X_{i})=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})$ (10)
where $\bm{h}^{i}_{j}$ and $X_{j}$ is the $j$-th row of $\bm{h}^{i}$ and $X$ respectively. $p^{k}_{\theta,i}$ is the mask probability in Equation 3. $\mathbb{H}$ is the entropy computation. The equality holds if and only if nodes are independent and hyperedges do not overlap. The second inequality is inspired by Kim et al. (2021). The proof of the proposition can be found in Appendix A.
Therefore, instead of optimizing the intractable objective 8, we optimize its upper bound (i.e. loss function) according to Equation 9 and 10.
$\mathcal{L}$ $=\|Y-\mathcal{F}_{l}\circ\mathcal{F}_{w}\circ\mathcal{F}_{c}(X)\|_{2}^{2}+\beta\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})\geq-I(H,Y)+\beta I(H,X)$ (11)
The learnable components are the shallow embeddings in Equation 3, the $\operatorname{DimReduction}$ MLP in Equation 6 and the LinearHead $\mathcal{F}_{l}$ in Equation 7. For how do we choose the trade-off coefficient $\beta$, see Appendix E.3 for more discussions.
4.3 Computational Complexity
Suppose we have $N$ regions as $N$ nodes and $K$ hyperedges. For the Constructor, the unignorable computation is from the mask operation. The computational complexity of this step for each hyperedge is $O(Nd)$ since it does a pairwise multiplication operation of a matrix of size $N\times d$. Given that there are $K$ hyperedges, the total complexity is $O(NKd)$. For the Weighter, the computation is from the dim reduction operation and the linear head. The dim reduction operation is an MLP. In this work, the hidden dimensions are a fraction of the original feature dimension. Therefore, the complexity of the dim-reduction MLP is $O(d^{2})$. The linear head only contributes $O(d)$, which is neglectable. As a result, the computational complexity of the whole model is $O(NKd+d^{2})=O(N^{2}K)$ since the feature dimension is equal to the number of regions (5.1). This complexity is just at the same scale as that of MLPs even though we are addressing a more challenging task: identifying high-order relationships in an exponential space.
@@ 5 Experiments
In this section, we conducted experiments to validate the quality of the learned hyperedges regarding the cognition phenotype outcome. Furthermore, we conducted ablation studies to validate the key components in our model. We also analyzed our results both quantitatively and qualitatively.
5.1 Experiment Settings
@@
To evaluate the quality of hyperedges obtained by HyBRiD, we use CPM (Shen et al., 2017), a standard model that could evaluate the correlation between connectivity and the prediction target, due to its high impact in the community. In the original implementation of CPM, weights of pairwise edges are obtained by Pearson correlation between nodes. These weights, as pairwise connectivity, are fed into the CPM. CPM will output a metric that measures the overall correlation between edges and the prediction target, which can be considered as a measure of edge qualities. This process is formulated as
$r^{\prime}=\operatorname{CPM}(\bm{w}_{p},Y)$ (12)
where $\bm{w}_{p}\in\mathbb{R}^{K_{p}}$ denotes the pairwise edge weights and $K_{p}$ is the total number of pairwise edges. $r^{\prime}$ is a metric that measures the quality of weights based on positive and negative correlations to the cognition score. To evaluate the quality of the learned weights for our model and baselines, we concatenate the learned weights $\bm{w}_{h}\in\mathbb{R}^{K_{h}}$ with the pairwise ones together into the CPM, and thus adjust Equation 12 to
$r=\operatorname{CPM}([\bm{w}_{p}\|\bm{w}_{h}],Y)$ (13)
@@
Datasets
We use functional brain imaging data from the first release of the Adolescent Brain Cognitive Development (ABCD) study, collected from $11,875$ children aged between 9 to 10 years old (Casey et al., 2018). The functional MRI (fMRI) data is collected from children when they were resting and when they performed three emotional and cognitive tasks. The fMRI data is processed using BioImage Suite (Joshi et al., 2011) with the standard preprocessing procedures, such as slice time and motion correction, registration to the MNI template (see details in Greene et al. (2018) and Horien et al. (2019)). We delete scans with more than 0.10 mm mean frame-to-frame displacement. We use the ABCD imaging data collected from the baseline (release 2.0) as well as the 2-year follow-up (release 3.0). For each time point, we included children subjected to four conditions: the resting state where the brain is not engaged in any activity (Rest), the emotional n-back task (EN-back), the Stop Signal task (SST), and the Monetary Incentive Delay (MID) task. In conclusion, we obtain $8$ datasets from $2$ timepoints and $4$ conditions (Rest, SST, EN-back, MID). Statistics of each dataset are summarized in Appendix C.
For the prediction target, we consider fluid intelligence as our label. Fluid intelligence reflects the general mental ability and plays a fundamental role in various cognitive functions. Recent literature has seen success in predicting intelligence based on pairwise predictive modeling (Dubois et al., 2018a; b), and we aim to improve the prediction accuracy of the current methods.
Data preprocessing
The raw fMRI data of an instance is represented in four dimensions (3 spatial dimensions + 1 temporal dimension), which can be imagined as a temporal sequence of 3D images. First, brain images are parceled into regions (or nodes) using the AAL3v1 atlas (Rolls et al., 2020). Following previous works (Kan et al., 2022b; Li et al., 2021; Thomas et al., 2022), each region’s time series is obtained by averaging all voxels in that region. Consistent with previous connectivity-based methods (Li et al., 2021; Kan et al., 2022b; Ktena et al., 2018; Said et al., 2023), for each region, we use its Pearson correlation coefficients to all regions as its features. We randomly split the data into train, validation, and test sets in a stratified fashion. The split ratio is 8:1:1.
Baselines
We compare our method with $3$ classes of baselines: 1) standard method, which is exactly the classical method that predicts outcomes based on pairwise edges (Shen et al., 2017; Dadi et al., 2019; Wang et al., 2021). The comparison with standard methods shows whether the high-order connectivity has its advantage over the classical pairwise one or not. 2) hypergraph construction methods. We consider $k$NN (Huang et al., 2009), $l_{1}$ hypergraph (Wang et al., 2015), and $l_{2}$ hypergraph (Jin et al., 2019). 3) Connectivity-based cognition prediction methods, which are state-of-the-art predictive models based on brain connectivity. We consider BrainNetGNN (Mahmood et al., 2021), BrainGNN (Li et al., 2021), and BrainNetTF (Kan et al., 2022b). BrainGB (Cui et al., 2022a) is a study of different brain graph neural network designs and we include its best design as a baseline. Note that since these models are unable to identify hyperedge structures of brain regions, we input their last layer embeddings (each entry as a weight) into the CPM model. Note that our weights $\bm{w}$ are also last layer embeddings in HyBRiD.
Implementation & Training Details
Due to the data scarcity, training on individual datasets would result in serious overfitting. To mitigate this, we train our model as well as baselines using all eight datasets together. Details of the overfitting issue and our parameter-sharing strategy are further discussed in Appendix E.1. See Appendix D for other implementation and training details such as hyperparameter choices and software/hardware specifications.
5.2 Quality of Hyperedges
We report $r$ values by CPM in Table 1. As we can see, HyBRiD outperforms the state-of-the-art predictive models on $7$ datasets, with an average improvement of $12.1\%$. Moreover, HyBRiD outperforms all other deep learning baselines in efficiency, with $87\%$ faster than the next fastest one (BrainNetTF). Refer to Appendix F for more runtime details.
@@ 5.2 Quality of Hyperedges
Table 1: $r$ values of our hyperedges compared to baselines. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
Ablation Studies
[…]
@@ 5.3 Further Analysis
We plot the hyperedge degree distribution in Fig.3(a). We find there are two distinct clusters in the figure. The first cluster is hyperedges with degree $\leq 5$. $1$-degree and $2$-degree hyperedges are special cases of our method: $1$-degree hyperedges are individual nodes, which imply the contribution of individual regions to the cognition. $2$-degree hyperedges reveal the importance of traditional pairwise connectivity. The other cluster concentrates around degree $25$, which implies the importance of relationships of multiple regions.
@@
CPM conducts a significance test on pairwise edges and hyperedges internally based on a linear regression model, and thus we can obtain a P-value for each hyperedge from the significance test. We define the significance of a hyperedge as $1-P_{v}\in[0,1]$ where $P_{v}$ is the P-value of that hyperedge.
Figure 4: Hyperedge profiles. (a) Hyperedge degree distribution of learned hyperedges. (b) Correlation between hyperedge degree and significance. (c) Comparison between the number hyperedges and pairwise edges under different significance thresholds. The total number of hyperedges is $32$. And the total number of pairwise edges is $26,896$.
The relationship between hyperedge degree and its significance is shown in Fig 3(b). In this figure, we find a strong positive correlation between a hyperedge’s degree and its significance, which indicates that interactions of multiple brain regions play more important roles in cognition than pairwise or individual ones. It is also worth mentioning that there is a turning point around degree $5$, which corresponds to the valley around $5$ in Figure 3(a).
@@
Figure 5: Visualization of the frequency of each region under different conditions.
To better understand the roles of each brain region in cognition under different conditions, we studied the frequency at which each region appears in a hyperedge out of all identified hyperedges. The frequency, which can be considered as a measure of region importance, is visualized in Figure 5. Visual regions (Fusiform, Cuneus, Calcarine) are especially active due to the intensive visual demands in all three conditions. We found that the Frontal_Inf_Oper and ACC_pre, recognized for their participation in response inhibition (Pornpattananangkul et al., 2016), frequently appear in the SST task. This aligns with what the SST task was designed to test. Interestingly, of the three conditions (SST, EN-back, MID), only EN-back prominently involves the Amygdala, a key region for emotion processing. This makes sense as EN-back is the only condition related to emotion. More visualizations and interpretations on the resting-state region importance and individual hyperedges can be found in Appendix G.

### 13 Feb 2024 v2

@@ 4.2 Optimization Framework
Since there are no existing IB frameworks that can be applied in our context, we propose a new IB framework named multi-head drop-bottleneck to optimize HyBRiD. To adopt an information bottleneck view of HyBRiD, we consider $X$, $Y$ and $H$ are random variables in the Markovian chain $X\leftrightarrow Y\leftrightarrow H$. According to our MIMR objective, we optimize
$\arg\max I(H;Y)-\beta I(H;X),$ (8)
where $I(\cdot;\cdot)$ denotes the mutual information. $I(H;Y)$ corresponds to the informativeness and $I(H;X)$ corresponds to the redundancy. $\beta$ is a coefficient trading off informativeness and redundancy. Since optimizing the mutual information for high-dimensional continuous variables is intractable, we instead optimize the lower bound of Equation 8. Specifically, for the first term (informativeness), it is easy to show
@@
$=\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]$
$\quad+\mathbb{E}_{p(H)}[\operatorname{KL}(p(Y|H)|q_{\phi}(Y|H))]$
$\geq\mathbb{H}[Y]+\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)],$
where $\mathbb{H}[\cdot]$ is the entropy computation. Since there is no learnable component in the entropy of $Y$, we only need to optimize the second term $\mathbb{E}_{p(Y,H)}[\log q_{\phi}(Y|H)]$. $q_{\phi}$ can be considered as a model that predicts $Y$ based on $H$, which essentially corresponds to $\mathcal{F}_{l}\circ\mathcal{F}_{w}$, where $\circ$ is the function composition. In practice, we set $q_{\phi}$ as a Gaussian model with variance $1$ as most probabilistic machine learning models do for continuous data modeling (Alemi et al., 2016; Luo et al., 2019; Peng et al., 2018; Kingma & Welling, 2013).
@@
$I(H;X)\leq\sum_{k=1}^{K}I(\boldsymbol{h}^{k};X)$ $\leq\sum_{k=1}^{K}\sum_{i=1}^{N}I(\boldsymbol{h}^{k}_{i};X_{i})$ (10)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i}),$
where $\boldsymbol{h}^{k}_{i}$ and $X_{i}$ is the $i$-th row of $\boldsymbol{h}^{k}$ and $X$ respectively. $p^{k}_{\theta,i}$ is the mask probability in Equation 3. $\mathbb{H}$ is the entropy computation. The equality holds if and only if nodes are independent and hyperedges do not overlap. It’s important to clarify that optimizing Equation 10 does not imply a penalty for overlaps. The second inequality is inspired by (Kim et al., 2021). The proof of the proposition can be found in Appendix A.
Therefore, instead of optimizing the intractable objective Equations 8, we optimize its upper bound (i.e. loss function) according to Equations 9 and 10.
$\mathcal{L}$ $=\|Y-\mathcal{F}_{l}\circ\mathcal{F}_{w}\circ\mathcal{F}_{c}(X)\|_{2}^{2}+\beta\sum_{k=1}^{K}\sum_{i=1}^{N}\mathbb{H}[X_{i}](1-p^{k}_{\theta,i})$ (11)
$\geq-I(H,Y)+\beta I(H,X).$
In conclusion, The learnable components are the shallow embeddings in Equation 3, the $\operatorname{DimReduction}$ MLP in Equation 6 and the LinearHead $\mathcal{F}_{l}$ in Equation 7. For how we choose the trade-off coefficient $\beta$, see Appendix H.2 for more discussions.
@@ 5 Experiments
In this section, we conduct experiments to validate the quality of the learned hyperedges in terms of the predictive performance towards the cognition phenotype outcome. Furthermore, we conduct ablation studies to validate the key components in our model. We also analyze our results both quantitatively and qualitatively.
5.1 Predictive Performance of Hypereges
Datasets
We consider two fMRI datasets:
1) Autism Brain Imaging Data Exchange (ABIDE) (Craddock et al., 2013) is an open-source dataset. This dataset involved resting-state fMRI of patients from 17 international sites, as well as the anatomical and phenotypic data. Regions are obtained by Craddock 200 atlas (Craddock et al., 2012). We use the preprocessed version from the official website. For prediction targets, we choose three intelligence quotients: FIQ (full-scale intelligence), VIQ (verbal intelligence quotient), and PIQ (performance intelligence).
2) Adolescent Brain Cognitive Development (ABCD) (Casey et al., 2018) is one of the largest public fMRI datasets. Access is limited and requires adherence to a rigorous data request procedure to acquire the data. The data is collected from $11,875$ children aged between 9 to 10 years old. The functional MRI (fMRI) data is collected from children when they were resting and when they performed three tasks (SST, EN-back, MID). We use the ABCD imaging data collected from the baseline (release 2.0) as well as the 2-year follow-up (release 3.0). In conclusion, we obtain $8$ sub-datasets (we refer to them as datasets from now on) from $2$ timepoints under $4$ tasks. Regions are obtained by AAL3v1 atlas (Rolls et al., 2020). For preprocess procedures of the ABCD dataset, please refer to Appendix D for more details. For the prediction target, we consider fluid intelligence as our label. Fluid intelligence reflects the general mental ability and plays a fundamental role in various cognitive functions.
For region features, consistent with previous connectivity-based methods (Li et al., 2021; Kan et al., 2022b; Ktena et al., 2018; Said et al., 2023), we use a region’s Pearson correlation coefficients to all other regions as the region features. Other details of the data preprocessing and statistics of each dataset are summarized in Appendix D.
@@ 5.1 Predictive Performance of Hypereges
To evaluate the quality of hyperedges obtained by HyBRiD, we use CPM (Shen et al., 2017), a standard model that could evaluate the relevance between the connectivity and the prediction target, due to its high impact in the community. In the original implementation of CPM, weights of pairwise edges are obtained by Pearson correlation between nodes. These weights, as pairwise connectivity, are fed into the CPM. CPM will output a metric that measures the overall correlation between edges and the prediction target, which can be considered as a measure of edge qualities. This process is formulated as
$r^{\prime}=\operatorname{CPM}(\boldsymbol{w}_{p},Y),$ (12)
where $\boldsymbol{w}_{p}\in\mathbb{R}^{K_{p}}$ denotes the pairwise edge weights and $K_{p}$ is the total number of pairwise edges. $r^{\prime}$ is a metric that measures the quality of weights based on positive and negative correlations to the outcome. To evaluate the quality of the learned weights for our model and baselines, we concatenate the learned weights $\boldsymbol{w}_{h}\in\mathbb{R}^{K_{h}}$ with the pairwise ones together into the CPM, and thus adjust Equation 12 to
$r=\operatorname{CPM}([\boldsymbol{w}_{p}\|\boldsymbol{w}_{h}],Y),$ (13)
@@
Type Model FIQ VIQ PIQ
Standard pairwise $0.052$ $0.124$ $0.056$
Hypergraph Construction $k$NN $0.023$ $0.093$ $0.056$
$l_{1}$ hypergraph $0.043$ $0.125$ $0.061$
$l_{2}$ hypergraph $0.148_{\pm 0.000}$ $0.141_{\pm 0.014}$ $0.063_{\pm 0.004}$
Connectivity based Prediction BrainNetGNN $\underline{0.162_{\pm 0.042}}$ $\underline{0.199_{\pm 0.042}}$ $\underline{0.223_{\pm 0.025}}$
BrainGB $0.125_{\pm 0.119}$ $0.154_{\pm 0.068}$ $0.157_{\pm 0.053}$
BrainGNN $0.105_{\pm 0.041}$ $0.176_{\pm 0.049}$ $0.159_{\pm 0.051}$
BrainNetTF $0.132_{\pm 0.111}$ $0.176_{\pm 0.053}$ $0.180_{\pm 0.054}$
Ours HyBRiD $\mathbf{0.181_{\pm 0.040}}$ $\mathbf{0.204_{\pm 0.031}}$ $\mathbf{0.245_{\pm 0.042}}$
Table 1: $r$ values of our hyperedges compared to baselines on the ABIDE dataset. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
@@
Table 2: $r$ values of our hyperedges compared to baselines on the ABCD dataset. Results are averaged over 10 runs. Deterministic methods do not have standard deviations.
Baselines
We compare our method with $3$ classes of baselines: 1) standard method, which is exactly the classical method that predicts outcomes based on pairwise edges (Shen et al., 2017; Dadi et al., 2019; Wang et al., 2021). The comparison with standard methods shows whether the high-order connectivity has its advantage over the classical pairwise one or not. 2) hypergraph construction methods. We consider $k$NN (Huang et al., 2009), $l_{1}$ hypergraph (Wang et al., 2015), and $l_{2}$ hypergraph (Jin et al., 2019). 3) connectivity-based phenotypic prediction methods, which are state-of-the-art predictive models based on brain connectivity. We consider BrainNetGNN (Mahmood et al., 2021), BrainGNN (Li et al., 2021), and BrainNetTF (Kan et al., 2022b). BrainGB (Cui et al., 2022a) is a study of different brain graph neural network designs and we include its best design as a baseline. Since none of these models are able to identify hyperedge structures of brain regions, we input their last layer embeddings (each entry as a weight) into the CPM model. Note that our weights $\boldsymbol{w}$ are also last layer embeddings in HyBRiD.
@@
Implementation & Training Details
Hyperparameter choices and other details can be found in Appendix E.
Results
We report $r$ values by CPM in Table 1 and Table 2. As we can see, on the ABIDE dataset, HyBRiD consistently outperforms all the baselines on different targets, with an average improvement of $8.8\%$ compared to the state-of-the-art model. On the ABCD dataset, HyBRiD outperforms the state-of-the-art predictive models on $7$ datasets of ABCD, with an average improvement of $12.1\%$. The results demonstrate our model is able to learn informative hyperedges towards different phenotypic outcomes from fMRI data of various brain states. Rest 1 is the only dataset that our model reaches the second-best. We attribute this imperfection to the discrepancy between validation and test datasets. This is because, through the inspection of the training/validation set performances, we find our model performs better than BrainNetTF on the training and validation dataset. More discussions about it can be found in Appendix F.
Runtime
HyBRiD outperforms all other deep learning baselines in efficiency, with $87\%$ faster than the second-fastest one (BrainNetTF). Refer to Appendix I for more runtime details.
Ablation Studies
[…]
Extra Experiments on Synthetic Dataset
Since there are no ground-truth hyperedges in real-world datasets of learning informative hyperedges towards a specific outcome, we construct a synthetic dataset to verify if our model can recover the correct hyperedge structure under the MIMR objective. We use the precision, recall, and F1 score to measure the correctness of the learned hyperedges with respect to the ground truth. Although it is challenging to learn hyperedges when only supervised by the task label, we find that our model reaches high performances, with an average improvement of 28.3% in terms of the F1 score, compared to the strongest baselines. Details about the synthetic experiments can be found in Appendix G.
Figure 4: Hyperedge profiles. (a) Hyperedge degree distribution of learned hyperedges. (b) Correlation between hyperedge degree and significance. (c) Comparison between the number of hyperedges and pairwise edges under different significance thresholds. The total number of hyperedges is $32$. And the total number of pairwise edges is $26,896$.
@@ 5.2 Further Analysis
In this subsection, we analyze the results of our model. We mainly use the ABCD dataset in the analysis since it is much larger than the ABIDE dataset.
@@
We plot the hyperedge degree distribution in Figure 3(a). We find there are two distinct clusters in the figure. The first cluster is hyperedges with degree $\leq 5$. $1$-degree and $2$-degree hyperedges are special cases of our method: $1$-degree hyperedges are individual nodes, which imply the contribution of individual regions to the cognition. $2$-degree hyperedges reveal the importance of traditional pairwise connectivity. The other cluster concentrates around degree $25$, which implies the importance of relationships of multiple regions.
@@
Since CPM conducts a significance test on pairwise edges and hyperedges internally based on a linear regression model, we can obtain a P-value for each hyperedge from the significance test. We define the significance of a hyperedge as $1-P_{v}\in[0,1]$ where $P_{v}$ is the P-value of that hyperedge.
The relationship between hyperedge degree and its significance is shown in Figure  3(b). In this figure, we find a strong positive correlation between a hyperedge’s degree and its significance, which indicates that interactions of multiple brain regions play more important roles in cognition than pairwise or individual ones. It is also worth mentioning that there is a turning point around degree $5$, which corresponds to the valley around $5$ in Figure 3(a).
@@
Figure 5: Visualization of the frequency of each region under different fMRI tasks.
To better understand the roles of each brain region in cognition under different fMRI tasks, we study the frequency at which each region appears in a hyperedge out of all identified hyperedges. The frequency, which can be considered as a measure of region importance, is visualized in Figure 5. Visual regions (Fusiform, Cuneus, Calcarine) are especially active due to the intensive visual demands in all three tasks. We found that the Inferior frontal gyrus, opercular part and the Anterior cingulate cortex, pregenual, recognized for their participation in response inhibition (Pornpattananangkul et al., 2016), frequently appear in the SST task. This aligns with what the SST task was designed to test. Interestingly, of the three tasks (SST, EN-back, MID), only EN-back prominently involves the Amygdala, a key region for emotion processing. This makes sense as EN-back is the only task related to emotion.
More visualizations of individual hyperedges can be found in Appendix J.